OpenStudy (anonymous):

Algebra 2 Will give MEDAL!

OpenStudy (anonymous):

What is the solution to the equation 4^2x ≈ 3? x = –0.631 x = –0.396 x = 0.396 x = 0.631

OpenStudy (anonymous):

@thomaster @Teinis

OpenStudy (anonymous):

@dumbsearch2

OpenStudy (anonymous):

Does it look like this? \[4^{2x} \approx 3\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, start by taking the natural log of both sides. This will allow us to drop down the 2x in front like this: \[2x \ln4 \approx \ln 3\]Do you see how we can isolate x now?

OpenStudy (anonymous):

By dividing it?

OpenStudy (anonymous):

We'll start by dividing both sides by ln 4, right. Now we'll just have 2x on the left: \[2x=\frac{ \ln 3 }{ \ln 4 }\]

OpenStudy (anonymous):

ok now to get x by itself we would need to divide both sides by 2 right?

OpenStudy (anonymous):

3/16

OpenStudy (anonymous):

@Juliette32801 that isn't in the answer choices though :(

OpenStudy (anonymous):

@AwesomeB I know hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3^(2x)=4 2xln(3)=ln(4) 2x=ln(4)/ln(3) 2x=1.26 x=.63092

OpenStudy (anonymous):

\[\approx \] 0.631 d.

OpenStudy (anonymous):

ty